OpenStudy (unklerhaukus):

Write down the electric field due to a point charge in cartesian co-ordinates and hence show that the curl of this field is zero.

OpenStudy (unklerhaukus):

needs to be in cartesian co-ordinates

OpenStudy (unklerhaukus):

\[r=\sqrt{x^2+y^2+z^2}\]\[\mathbf r = x\,\hat{\mathbf x}+y\,\hat{\mathbf y}+z\,\hat{\mathbf z}\]

OpenStudy (unklerhaukus):

\[\begin{align}\mathbf E &= K\frac{Q}{r^2}\hat{\mathbf r}\\ &=K\frac{Q}{r^3}{\mathbf r}\\ &=K\frac Q{(x^2+y^2+z^2)^{3/2}}(x\,\hat{\mathbf x}+y\,\hat{\mathbf y}+z\,\hat{\mathbf z}) \end{align}\]

OpenStudy (unklerhaukus):

\[ E_x = \frac{Qx}{(x^2+y^2+z^2)^{3/2}}\\ E_y = \frac{Qy}{(x^2+y^2+z^2)^{3/2}}\\ E_z = \frac{Qz}{(x^2+y^2+z^2)^{3/2}} \]

OpenStudy (unklerhaukus):

\[\nabla \times\mathbf E = \begin{vmatrix}\hat{\mathbf x}&\hat{\mathbf y}&\hat{\mathbf z}\\\frac\partial{\partial x}&\frac\partial{\partial y}&\frac\partial{\partial z}\\E_x&E_y&E_z \end{vmatrix}\\ \qquad \quad=\left(\frac{\partial E_z}{\partial y}-\frac{\partial E_y}{\partial z}\right)\hat{\mathbf x} +\left(\frac{\partial E_x}{\partial z}-\frac{\partial E_z}{\partial x}\right)\hat{\mathbf y} +\left(\frac{\partial E_y}{\partial x}-\frac{\partial E_x}{\partial y}\right)\hat{\mathbf x}\]

OpenStudy (unklerhaukus):

\[E_x = \frac{Qx}{(x^2+y^2+z^2)^{3/2}}\] \[\frac{\partial E_x}{\partial y} = \]

OpenStudy (unklerhaukus):

i think i have got it , but the working is horrible

OpenStudy (unklerhaukus):

Is there a simpler method?

OpenStudy (sidsiddhartha):

so we have to prove that the field is irrotational right?

OpenStudy (sidsiddhartha):

we can assume the field \[E=E_x \hat{x}+E_y \hat{y}+E_z \hat{z}\]

OpenStudy (sidsiddhartha):

and corresponding potential field \[V=V_x \hat{x}+V_y \hat{y}+V_z \hat{z}\]

OpenStudy (sidsiddhartha):

now we can use the relation \[E=- \nabla V\]

OpenStudy (sidsiddhartha):

it will produce \[\nabla \times E=- \nabla \times (\nabla V)\]

OpenStudy (sidsiddhartha):

makes sense?

OpenStudy (sidsiddhartha):

now using this relation -